# Lorentz-violating extension of Wigner function formalism and chiral kinetic theory

### v1

@@ Abstract
The quantum kinetic equation for the gauge-invariant Wigner function, constructed from spinor fields that obey the Dirac equation modified by CPT and Lorentz symmetry-violating terms, is presented. The equations for the components of Wigner function in the Clifford algebra basis are accomplished. Focusing on the massless case, an extended semiclassical chiral kinetic theory in the presence of external electromagnetic fields is developed. We calculate the chiral currents and establish the anomalous magnetic and separation effects in a Lorentz-violating background. The chiral anomaly within the context of extended Quantum Electrodynamics is elucidated. Finally, we derive the semiclassical Lorentz-violating extended chiral transport equation.
@@ I Introduction
The fermionic constituents of the Standard Model of particle physics are mainly the charged Dirac and Weyl particles which also appear in condensed matter as quasiparticles. Various methods have been employed to study the dynamical features of spin-1/2 fermions. Nevertheless, an intuitive understanding of physical phenomena involving them is offered by kinetic theory. A systematic construction of kinetic theory begins with the quantum kinetic equation satisfied by the relativistic Wigner function [1, 2]. It is the relativistic quantum field theory counterpart of the classical distribution function. In high energy physics, quantum kinetic theory has predominately been investigated to elucidate phenomena arising in the heavy-ion collisions due to the chiral nature of quarks in a novel phase of quark-gluon plasma. A a comprehensive review of quantum transport theory derived by means of Wigner function method and its applications, is provided in [3].
@@
One of the main properties of quantum field theory models is the invariance under Lorentz transformation. However, there are debates about the violation of Lorentz symmetry under certain conditions, such as very high energies. For instance, it may arise from the String Theory as outlined in [4, 5]. On the other hand, Lorentz symmetry violation may stem from anisotropy in spacetime. A brief historical overview and current status of ideas regarding potential sources of Lorentz violation in particle physics can be found in [6]. Our discussion is based on the Standard Model Extension (SME), an effective field theory that incorporates Lorentz and CPT violating terms into the standard model of particle physics [7, 8]. The SME assumes that a fundamental theory, which may be String Theory, undergoes spontaneous Lorentz symmetry breaking. This results in a low-energy effective action with explicit Lorentz symmetry-violating (LSV) terms while maintaining microcausality, positivity of energy, and energy-momentum conservation. Notably, in SME particle and observer Lorentz transformations are not treated on equal footing. Observer frames differ in orientation and velocity, hence they are related to coordinate changes. The fundamental theory’s Lorentz symmetry is spontaneously broken, so that observer Lorentz symmetry is conserved. However, particle Lorentz symmetry is broken because it is defined by rotations and boosts of the localized field while keeping the expectation values of tensor fields unchanged. Restricting the minimal SME to Abelian gauge theory one obtains the extended Quantum Electrodynamics (QED) whose fermionic sector is given by the Dirac spinors coupled to electromagnetic gauge fields in the presence of Lorentz and CPT violating terms. This framework is also utilized to discuss violation of emergent Lorentz symmetry in Dirac and Weyl semimetals [9].
@@
We deal with the modified Dirac equation with a certain set of LSV terms: $a_{\mu}$ and $c_{\mu\nu}.$ The Boltzmann equation for this choice of LSV coefficients has been examined in [10]. The spinor fields obeying the modified Dirac equation can be quantized using the customary methods of field theory, and the extended QED is invariant under the gauge transformations. Thus, one can construct the extended relativistic gauge-invariant Wigner function and derive the quantum kinetic equation as in the ordinary relativistic kinetic theory. We expand the Wigner function in Clifford algebra basis where the coefficients are scalar, pseudoscalar, vector, axial-vector and tensor field variables. The quantum kinetic equation yields the coupled equations of these fields, which simplify when the mass is set to zero. We then consider the massless particles and explore the chiral kinetic theory which is manifestly observer Lorentz invariant. One of the peculiar aspects of chiral fermions is the emergence of anomalous effects like chiral magnetic [11, 12] and separation [13, 14, 15] effects. By choosing the distribution function appropriately we will be able to calculate vector and axial-vector particle currents and show that these anomalous effects get contribution from the LSV terms. Moreover, by studying their conservation we will demonstrate that the chiral anomaly is also altered.
@@
In the next section we discuss the relativistic gauge-invariant Wigner function and the quantum kinetic equation for Dirac particles within extended QED. In Section III the quantum kinetic equations of the chiral vector fields will be presented. Two of the three coupled equations are solved in a semiclassical approach. In Section IV we employ the modified Fermi-Dirac distribution function [10] and calculate the chiral vector fields, which are then used to establish four-currents generating anomalous magnetic and separation effects. The vector current is shown to be conserved, while the axial-vector four-current leads to the chiral anomaly. In Section V, an observer Lorentz invariant semiclassical chiral kinetic equation is derived. It is followed by a discussion of results and potential future studies in the final section.
@@ II Quantum kinetic equation
We deal with the Dirac fermion of mass $m,$ and charge $q,$ in the presence of external electromagnetic fields. There may be several Lorentz symmetry violating terms that are coordinate reparametrization and gauge invariant [16]. However, we consider a restricted set of Lorentz symmetry violating coefficients described by the Lagrangian density of the Dirac spinors $\psi$ coupled to the electromagnetic gauge field $A_{\mu},$ given as $(c=1)$
@@
The Minkowski metric is $g_{\mu\nu}={\rm diag}(1,-1,-1,-1),$ and $\gamma_{\mu}$ are the ordinary gamma matrices satisfying $\{\gamma_{\mu},\gamma_{\nu}\}=2g_{\mu\nu}.$ The LSV coefficients $a_{\mu}$ and $c_{\mu\nu}$ are real and constant. The former violates also the $CPT$ invariance. By varying (1) with respect to $\bar{\psi}$ and $\psi,$ one derives the equations of motion as
@@
In [7] and [16] it has been demonstrated that the spinor operators can be defined as in the ordinary case: In the “concordant frame” where the LSV coefficients are small, there exists a transformation of the spinors $\psi(x)$ that leads to a free Hamiltonian possessing two positive and two negative eigenvalues. Hence, one introduces plane wave solutions and define wave packets. Then, one proceeds as in the ordinary quantum field theory and introduce the spinor operators $\hat{\psi}(x)$ and $\hat{\bar{\psi}}(x).$ Therefore, by means of the four-momentum $p^{\mu}$ one can define the Wigner operator as
@@
Let us choose $c_{\mu\nu}=c_{\nu\mu},$ without loss of generality, and introduce
@@
In the sequel we consider the chiral fermions in the semiclassical limit.
@@ III Chiral vector fields
In fact, we will deal with the vector field $\mathcal{V}_{\mu}$ and the axial-vector field $\mathcal{A}_{\mu}$ for $m=0$ in the semiclassical limit, keeping at most the $\hbar$ dependent terms. Thus, in terms of
@@
where $\chi=\pm$ labels the right-handed and left-handed vector fields. Then (20)-(25) can be combined into
@@
where $f_{\chi}^{(0)}$ is a general distribution function. (30) should also satisfy (28):
@@
Now, we plug (42) into (37) and multiply it with $\bar{p}^{\nu}:$
@@
By means of the Schouten identity (38) and the modified Vlasov equation (34), we can show that it indeed satisfies (46):
@@
where we defined the distribution function $f_{\chi}=f_{\chi}^{(0)}+\hbar f_{\chi}^{(1)}.$ Obviously, it should also satisfy (28). We will come back to it in Section V. Now, we will specify the equilibrium distribution function and calculate chiral currents.
@@ IV Fermi-Dirac distribution and the chiral currents
The equilibrium distribution functions of the particles which obey Bose-Einstein and Fermi-Dirac statistics in the LSV background are studied in [10]. For massless fermions it is given by
@@
We set the Boltzmann constant $k=1.$ In terms of the four-velocity of the fluid $u_{\mu},$ we defined
$\tilde{u}_{\mu}=G^{-1}_{\mu\nu}u^{\nu},$ (50)
which satisfies
$\tilde{u}_{\mu}\tilde{u}^{\mu}=1.$ (51)
$\mu_{\chi}$ are given by the total chemical potential $\mu$ and the chiral chemical potential $\mu_{5}$ as $\mu_{{\scriptscriptstyle{R}},{\scriptscriptstyle{L}}}=\mu\pm\mu_{5}.$ We would like to study the chiral vector field (48) by choosing $f_{\chi}$ as the modified Fermi-Dirac distribution function (49) and setting $n=\tilde{u}.$ In this frame we introduce $\bar{E}_{\mu}=\bar{F}_{\mu\nu}\tilde{u}^{\nu},$ and $\bar{B}_{\mu}=(1/2)\epsilon_{\mu\nu\alpha\beta}\tilde{u}^{\nu}\bar{F}^{\alpha\beta},$ so that the field strength and its dual can be expressed as
@@
Let us deal with vanishing vorticity $\partial^{\beta}u^{\kappa}=0,$ and constant electromagnetic fields, $E_{\mu},B_{\mu}.$ Then, by employing the identity $\bar{p}^{2}\delta^{\prime\prime}(\bar{p}^{2})=-2\delta^{\prime}(\bar{p}^{2})$ and the relation
@@
$\xi_{\chi}$ $=$ $\frac{}{4\pi^{2}\hbar^{2}|G|}\left(\mu_{\chi}+\tilde{u}\cdot a\right).$ (62)
Then, the vector and axial-vector currents defined by $j^{\mu}=j^{\mu}_{R}+j^{\mu}_{L}$ and $j^{\mu}_{5}=j^{\mu}_{R}-j^{\mu}_{L},$ are established as
@@
In the absence of LSV terms, the magnetic parts of the currents (63) and (64) are known as chiral magnetic and separation effects. We found that these effects are modified in two ways. First, the magnetic field is replaced by $\bar{B}$, whose dependence on the magnetic field $B_{\mu}=(1/2)\epsilon_{\mu\nu\alpha\beta}u^{\nu}F^{\alpha\beta},$ is not direct. Second, the coefficients (67), (68) depend on the LSV terms.
@@
The chiral anomaly in this context has been studied in [23] and [24]. In [23] $a=0,$ and the chiral current was defined as $G_{\mu\nu}J^{\nu}_{5}$ where $J_{5}$ is the ordinary axial-vector current. Thus, its magnetic part is similar to (63). By studying the index theorem, they concluded that the chiral anomaly is the same as in the ordinary Lorentz invariant theory. However, in [24] chiral anomaly was calculated by employing the Fujikawa method as
@@ V Chiral transport equation
Following the approach in [21], we will show that (71) leads to the LSV-extended chiral transport equation. Since $\bar{F}_{\mu\nu}$ is antisymmetric, the first term can easily be shown to be
@@
Obviously, one should keep only the terms up to the $\hbar$ order. (78) is the semiclassical chiral transport equation, which is observer Lorentz invariant. The delta function dictates the mass shell condition.

### v2

@@ Abstract
The quantum kinetic equation for the gauge-invariant Wigner function, constructed from spinor fields that obey the Dirac equation modified by $CPT$ and Lorentz symmetry-violating terms, is presented. The equations for the components of the Wigner function in the Clifford algebra basis are accomplished. Focusing on the massless case, an extended semiclassical chiral kinetic theory in the presence of external electromagnetic fields is developed. We calculate the chiral currents and establish the anomalous magnetic and separation effects in a Lorentz-violating background. The chiral anomaly within the context of extended quantum electrodynamics is elucidated. Finally, we derive the semiclassical Lorentz-violating extended chiral transport equation.
@@ I Introduction
The fermionic constituents of the Standard Model of particle physics are mainly the charged Dirac and Weyl particles which also appear in condensed matter as quasiparticles. Various methods have been employed to study the dynamical features of spin-1/2 fermions. Nevertheless, an intuitive understanding of physical phenomena involving them is offered by kinetic theory. A systematic construction of kinetic theory begins with the quantum kinetic equation satisfied by the relativistic Wigner function [1, 2]. It is the relativistic quantum field theory counterpart of the classical distribution function. In high-energy physics, quantum kinetic theory has predominately been investigated to elucidate phenomena arising in the heavy-ion collisions due to the chiral nature of quarks in a novel phase of quark-gluon plasma. A comprehensive review of quantum transport theory derived by means of the Wigner function method and its applications is provided in [3].
@@
One of the main properties of quantum field theory models is the invariance under Lorentz transformation. However, there are debates about the violation of Lorentz symmetry under certain conditions, such as very high energies. For instance, it may arise from the string theory as outlined in [4, 5]. On the other hand, Lorentz symmetry violation may stem from anisotropy in spacetime. A brief historical overview and current status of ideas regarding potential sources of Lorentz violation in particle physics can be found in [6]. Our discussion is based on the Standard Model extension (SME), an effective field theory that incorporates Lorentz- and $CPT$-violating terms into the Standard Model of particle physics [7, 8]. The SME assumes that a fundamental theory, which may be string theory, undergoes spontaneous Lorentz symmetry breaking. This results in a low-energy effective action with explicit Lorentz symmetry-violating (LSV) terms while maintaining microcausality, positivity of energy, and energy-momentum conservation. Notably, in SME particle and observer Lorentz transformations are not treated on equal footing. Observer frames differ in orientation and velocity, hence they are related to coordinate changes. The fundamental theory’s Lorentz symmetry is spontaneously broken, so that observer Lorentz symmetry is conserved. However, particle Lorentz symmetry is broken because it is defined by rotations and boosts of the localized field while keeping the expectation values of tensor fields unchanged. Restricting the minimal SME to Abelian gauge theory one obtains the extended quantum electrodynamics (QED) whose fermionic sector is given by the Dirac spinors coupled to electromagnetic gauge fields in the presence of Lorentz- and $CPT$-violating terms. This framework is also utilized to discuss violation of emergent Lorentz symmetry in Dirac and Weyl semimetals [9].
@@
We deal with the modified Dirac equation with a certain set of LSV terms: $a_{\mu}$ and $c_{\mu\nu}.$ The Boltzmann equation for this choice of LSV coefficients has been examined in [10]. The spinor fields obeying the modified Dirac equation can be quantized using the customary methods of field theory, and the extended QED is invariant under the gauge transformations. Thus, one can construct the extended relativistic gauge-invariant Wigner function and derive the quantum kinetic equation as in the ordinary relativistic kinetic theory. We expand the Wigner function in the Clifford algebra basis where the coefficients are scalar, pseudoscalar, vector, axial-vector, and tensor field variables. The quantum kinetic equation yields the coupled equations of these fields, which simplify when the mass is set to zero. We then consider the massless particles and explore the chiral kinetic theory which is manifestly observer Lorentz invariant. One of the peculiar aspects of chiral fermions is the emergence of anomalous effects like chiral magnetic [11, 12] and separation [13, 14, 15] effects. By choosing the distribution function appropriately we will be able to calculate vector and axial-vector particle currents and show that these anomalous effects get contribution from the LSV terms. Moreover, by studying their conservation we will demonstrate that the chiral anomaly is also altered.
@@
In the next section we discuss the relativistic gauge-invariant Wigner function and the quantum kinetic equation for Dirac particles within extended QED. In Sec. III the quantum kinetic equations of the chiral vector fields will be presented. Two of the three coupled equations are solved in a semiclassical approach. In Sec. IV we employ the modified Fermi-Dirac distribution function [10] and calculate the chiral vector fields, which are then used to establish four-currents generating anomalous magnetic and separation effects. The vector current is shown to be conserved, while the axial-vector four-current leads to the chiral anomaly. In Sec. V, an observer Lorentz-invariant semiclassical chiral kinetic equation is derived. It is followed by a discussion of results and potential future studies in the final section.
@@ II Quantum kinetic equation
We deal with the Dirac fermion of mass $m$ and charge $q$ in the presence of external electromagnetic fields. There may be several Lorentz symmetry-violating terms that are coordinate reparametrization and gauge invariant [16]. However, we consider a restricted set of Lorentz symmetry-violating coefficients described by the Lagrangian density of the Dirac spinors $\psi$ coupled to the electromagnetic gauge field $A_{\mu},$ given as $(c=1)$
@@
The Minkowski metric is $g_{\mu\nu}={\rm diag}(1,-1,-1,-1),$ and $\gamma_{\mu}$ are the ordinary $\gamma$ matrices satisfying $\{\gamma_{\mu},\gamma_{\nu}\}=2g_{\mu\nu}.$ The LSV coefficients $a_{\mu}$ and $c_{\mu\nu}$ are real and constant. The former violates also the $CPT$ invariance. By varying (1) with respect to $\bar{\psi}$ and $\psi,$ one derives the equations of motion as
@@
In [7, 16] it has been demonstrated that the spinor operators can be defined as in the ordinary case: In the “concordant frame” where the LSV coefficients are small, there exists a transformation of the spinors $\psi(x)$ that leads to a free Hamiltonian possessing two positive and two negative eigenvalues. Hence, one introduces plane wave solutions and define wave packets. Then, one proceeds as in the ordinary quantum field theory and introduces the spinor operators $\hat{\psi}(x)$ and $\hat{\bar{\psi}}(x).$ Therefore, by means of the four-momentum $p^{\mu}$ one can define the Wigner operator as
@@
Let us choose $c_{\mu\nu}=c_{\nu\mu}$ without loss of generality and introduce
@@
In the subsequent sections we consider the chiral fermions in the semiclassical limit.
@@ III Chiral vector fields
In fact, we will deal with the vector field $\mathcal{V}_{\mu}$ and the axial-vector field $\mathcal{A}_{\mu}$ for $m=0$ in the semiclassical limit, keeping at most the $\hbar$-dependent terms. Thus, in terms of
@@
where $\chi=\pm$ labels the right- and left-handed vector fields. Then (20)-(25) can be combined into
@@
where $f_{\chi}^{(0)}$ is a general distribution function. Equation (30) should also satisfy (28),
@@
Now, we plug (42) into (37) and multiply it with $\bar{p}^{\nu},$
@@
By means of the Schouten identity (38) and the modified Vlasov equation (34), we can show that it indeed satisfies (46),
@@
where we defined the distribution function $f_{\chi}=f_{\chi}^{(0)}+\hbar f_{\chi}^{(1)}.$ Obviously, it should also satisfy (28). We will come back to it in Sec. V. Now, we will specify the equilibrium distribution function and calculate chiral currents.
@@ IV Fermi-Dirac distribution and the chiral currents
The equilibrium distribution function of the particles which obey Fermi-Dirac statistics in the LSV background is studied in [10]. There, the standard relativistic kinetic theory formulation of gases [22] is employed by incorporating the LSV modifications of the equations of motion: The massless particles obey the extended dispersion relation
$\bar{p}\cdot\bar{p}=0,$ (49)
and the constraint
$G^{-1}_{\rho\mu}u^{\mu}\left(G^{-1}\right)^{\rho\nu}u_{\nu}-1=0.$ (50)
$u_{\mu}\equiv dx_{\mu}/d\tau$ is the four-velocity of the fluid where $\tau$ is the proper time in the absence of LSV. By introducing
$\tilde{u}_{\mu}\equiv G^{-1}_{\mu\nu}u^{\nu},$ (51)
(50) can be expressed as
$\tilde{u}_{\mu}\tilde{u}^{\mu}=1.$ (52)
The Boltzmann and relativistic Uehling-Uhlenbeck equations have been derived following the ordinary relativistic theory in terms of the one-particle phase space distribution function $f(x,p),$ and $\tilde{p}^{\mu}=G^{\mu\nu}\bar{p}_{\nu}.$ Then, the H-theorem is demonstrated by making use of the extended transport equations and introducing the entropy-density four-current
$s^{\mu}=\frac{1}{(2\pi\hbar)^{3}}\int\frac{d^{3}p}{\tilde{p}^{0}}\tilde{p}^{\mu}\left[\left(f(x,p)-1\right)\ln\left(1-f(x,p)\right)-f(x,p)\ln f(x,p)\right],$ (53)
which yields the total entropy
$S=\int d^{3}xs^{0}=\frac{1}{(2\pi\hbar)^{3}}\int d^{3}xd^{3}p\,\left[\left(f(x,p)-1\right)\ln\left(1-f(x,p)\right)-f(x,p)\ln f(x,p)\right].$ (54)
The Boltzmann constant is set $k=1.$ Observe that (54) is independent of the LSV coefficients. Thus, as in the ordinary case one can observe that for particles with momentum $p_{1},p_{2}$ scattered to particles with momentum $p_{3},p_{4},$ the total entropy (54) is stationary when $\phi(x,p)=-\ln\left[f(x,p)/(1-f(x,p))\right]$ satisfies the condition
$\phi(x,p_{3})+\phi(x,p_{4})-\phi(x,p_{1})-\phi(x,p_{2})=0.$ (55)
For momentum conserving scatterings, the most general solution of (55) can be shown to be
$f(x,p)=\frac{1}{e^{-\alpha(x)+\beta(x)\cdot p}+1},$ (56)
where $\alpha(x)$ and $\beta_{\mu}(x)$ are arbitrary. Equation (56) should obey the Uehling-Uhlenbeck equation which is solved by the Fermi-Dirac distribution function
$f_{FD}(x,p)=\frac{1}{e^{(-\mu+u\cdot p)/T}+1},$ (57)
for $\partial u_{\mu}/\partial x^{\nu}=0.$ On mass-shell $\bar{p}^{2}=0,$ thus the equilibrium distribution function for chiral fermions and antifermions is given by
$f_{\chi}^{eq}$ $=$ $\frac{2}{(2\pi\hbar)^{3}}\left[\frac{\theta(\tilde{u}\cdot\bar{p})}{e^{(u\cdot p-\mu_{\chi})/T}+1}+\frac{\theta(-\tilde{u}\cdot\bar{p})}{e^{-(u\cdot p-\mu_{\chi})/T}+1}\right].$ (58)
Now by expressing it in $\bar{p}$ we get
@@
$\mu_{\chi}$ are given by the total chemical potential $\mu$ and the chiral chemical potential $\mu_{5}$ as $\mu_{{\scriptscriptstyle{R}},{\scriptscriptstyle{L}}}=\mu\pm\mu_{5}.$ Observe that the chemical potentials $\mu_{\chi}$ are effectively shifted with $\tilde{u}\cdot a,$ reminiscent of expressing the distribution function (58) in terms of $\bar{p}_{\mu},$ which is the LSV extended momentum appearing in the dispersion relation (49).
We would like to study the chiral vector field (48) by choosing $f_{\chi}$ as the modified Fermi-Dirac distribution function (59) and setting $n=\tilde{u}.$ In this frame we introduce $\bar{E}_{\mu}=\bar{F}_{\mu\nu}\tilde{u}^{\nu},$ and $\bar{B}_{\mu}=(1/2)\epsilon_{\mu\nu\alpha\beta}\tilde{u}^{\nu}\bar{F}^{\alpha\beta},$ so that the field strength and its dual can be expressed as
@@
Let us deal with vanishing vorticity $\partial^{\beta}u^{\kappa}=0$ and constant electromagnetic fields $E_{\mu},B_{\mu}.$ Then, by employing the identity $\bar{p}^{2}\delta^{\prime\prime}(\bar{p}^{2})=-2\delta^{\prime}(\bar{p}^{2})$ and the relation
@@
$\xi_{\chi}$ $=$ $\frac{1}{4\pi^{2}\hbar^{2}|G|}\left(\mu_{\chi}+\tilde{u}\cdot a\right).$ (70)
Then, the vector and axial-vector currents defined by $j^{\mu}=j^{\mu}_{R}+j^{\mu}_{L}$ and $j^{\mu}_{5}=j^{\mu}_{R}-j^{\mu}_{L}$ are established as
@@
In the absence of LSV terms, the magnetic parts of the currents (71) and (72) are known as chiral magnetic effect (CME) and chiral separation effect (CSE). We found that these effects are modified in two ways. First, the magnetic field is replaced by $\bar{B}$, whose dependence on the magnetic field $B_{\mu}=(1/2)\epsilon_{\mu\nu\alpha\beta}u^{\nu}F^{\alpha\beta}$ is not direct. Second, the coefficients (75) and (76) depend on the LSV terms.
@@
The chiral anomaly in this context has been studied in [24, 25]. In [24] $a=0,$ and the chiral current was defined as $G_{\mu\nu}J^{\nu}_{5}$ where $J_{5}$ is the ordinary axial-vector current. Thus, its magnetic part is similar to (71). By studying the index theorem, they concluded that the chiral anomaly is the same as in the ordinary Lorentz-invariant theory. The same conclusion is drawn in [26]. However, in [25], the chiral anomaly was calculated by employing the Fujikawa method as
@@ V Chiral transport equation
Following the approach in [21], we will show that (79) leads to the LSV extended chiral transport equation. Since $\bar{F}_{\mu\nu}$ is antisymmetric, the first term can easily be shown to be
@@
Obviously, one should keep only the terms up to the $\hbar$ order. Equation (86) is the semiclassical chiral transport equation, which is observer Lorentz invariant. The $\delta$ function dictates the mass-shell condition.